OpenStudy (anonymous):

Solve the trig equation. Apparently the answer is No solution because cos cant be more than 2:

OpenStudy (anonymous):

\[\huge \cos^{2}x + \cos(2x) = \frac{5}{4}\]

OpenStudy (anonymous):

I can't seem to get No Solution though...

OpenStudy (istim):

Anyone going to help? I got a 60% in that class.

OpenStudy (anonymous):

Hey @saifoo.khan any ideas?

OpenStudy (saifoo.khan):

@Study23 , access is awesome with this!

OpenStudy (accessdenied):

There should be a solution to this one. Try using a trig identity for cos(2x) like 2cos^2 x - 1

OpenStudy (anonymous):

@AccessDenied I did. But it didn't work...

OpenStudy (accessdenied):

Could you post your process?

OpenStudy (anonymous):

Move the 5/4 over to the left and make it cos^2x+cos(2x)-5/4=0 and you can use quadratic equation after that

OpenStudy (accessdenied):

Sorry, was hoping we could slip in just before the break! I used the identity \(cos(2x) = cos^2 (x) - 1\). The rest of the process is mostly Algebra: getting the \(cos^2 x\) isolated, taking the positive and negative square root of both sides, and then using the inverse to get \(x\). You can't use it in the form of a quadratic because the linear term is a \(cos(2x)\) and the leading term is a \((cos(x))^2\).

OpenStudy (anonymous):

@AccessDenied I used the Quadratic formula... so that isnt correct?

OpenStudy (accessdenied):

oops i posted the identity incorrectly: \(cos(2x) = 2cos^2 x - 1\) but yeah, quadratic formula only works if you have \(cos(x)\) for the terms

OpenStudy (campbell_st):

use cos(2x) = 2cos^(x) - 1 so the problem is now cos^2(x) + 2cos^2(x) - 1 = 5/4 3cos^2(x) - 1 = 5/4 3cos^2(x) = 9/4 cos^2(x) = 3/4 \[\cos(x) = \pm \sqrt{3}/2\] there are solutions to this problem in each quadrant I'll assume you are working in radians \[\cos^{-1}(\sqrt{3}/2) = pi /3\]